OpenStudy (anonymous):

a single die is rolled one time. Find the probability of rolling a greater than 5 or less than 2

OpenStudy (anonymous):

so the event space has only {1,6} out of {1,2,3,4,5,6} so reqd prob =2/6=1/3

OpenStudy (mosaic):

A die has 6 possible outcomes: 1, 2, 3, 4, 5, 6 Here we are interested in greater than 5, which means 6, and less than 2, which means 1. So interested outcome is: 1 or 6 2 out of 6 or 1/3

OpenStudy (anonymous):

I had 1/3 just wanted to be sure...thx